OpenStudy (anonymous):

If f(x) = 2x - 9, then f(3) = -3. True False

hartnn (hartnn):

f(3)=2(3)-9 = 6- 9 = -3 so what do u think ?

OpenStudy (anonymous):

true or false

OpenStudy (jwheele1):

f(x) means the same thin as Y

OpenStudy (jwheele1):

so what you see then is that y = -3

OpenStudy (jwheele1):

if y = -3 than you can set it up this way: -3 = 2x - 9

OpenStudy (jwheele1):

then you just need to solve for x. Can you do that part?

OpenStudy (anonymous):

TRUE OR FALSE!

OpenStudy (jwheele1):

hey man...that just simply is not how this site works

OpenStudy (anonymous):

WHY NOT?

OpenStudy (anonymous):

\[f(x) = 2x - 9\] \[f(3) = 2*3 - 9 = -3\]

OpenStudy (anonymous):

So True..

OpenStudy (jwheele1):

this is a place to come and learn how to do things with help not get the answers

hartnn (hartnn):

because this is not an answering site!

hartnn (hartnn):

this is learning site

hartnn (hartnn):

u come here to learn, not to get answers

OpenStudy (anonymous):

sorry for caps

OpenStudy (jwheele1):

I mean, I am literally sitting at a computer at college during my free time to help students out because I want to help. I am not interested in wasting my time doing someones homework for them. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ok ok gosh

OpenStudy (jwheele1):

Its already been solved for you so I am just gonna work out the steps at this point. -3 = 2x - 9 goal: to get the number attached to the variable by itself. to do this you must add 9 to both sides 6 = 2x Goal: t get the variable all by itself. To do this you must divide both sides by the number glued to the variable. which is 2. 6/2 = 3 2/2 = 1 so, 3 = 1x or x = 3

OpenStudy (jwheele1):

which is shown by f(3)= -3.